### v1

# ToonifyGB: StyleGAN-based Gaussian Blendshapes for 3D Stylized Head Avatars

Abstract.
The introduction of 3D Gaussian blendshapes has enabled the real-time reconstruction of animatable head avatars from monocular video. Toonify, a StyleGAN-based framework, has become widely used for facial image stylization. To extend Toonify for synthesizing diverse stylized 3D head avatars using Gaussian blendshapes, we propose an efficient two-stage framework, ToonifyGB. In Stage 1 (stylized video generation), we employ an improved StyleGAN to generate the stylized video from the input video frames, which addresses the limitation of cropping aligned faces at a fixed resolution as preprocessing for normal StyleGAN. This process provides a more stable video, which enables Gaussian blendshapes to better capture the high-frequency details of the video frames, and efficiently generate high-quality animation in the next stage. In Stage 2 (Gaussian blendshapes synthesis), we learn a stylized neutral head model and a set of expression blendshapes from the generated video. By combining the neutral head model with expression blendshapes, ToonifyGB can efficiently render stylized avatars with arbitrary expressions. We validate the effectiveness of ToonifyGB on the benchmark dataset using two styles: Arcane and Pixar.
Toonify, StyleGAN, Gaussian Splatting, Gaussian Blendshapes, 3D Stylized Head Avatar, Facial Animation
journal: TOG
1. Introduction
With the advancement of 3D head reconstruction technologies, individuals are now able to personalize unique avatars for telepresence and virtual/augmented reality (VR/AR) applications, which serve as an essential foundation for the rise of the metaverse. Considering user preferences and privacy concerns, the creation of stylized avatars is an important topic that merits further investigation. Toonify (Pinkney and Adler, 2020) is a StyleGAN‑based method for stylizing 2D facial images. In contrast to photo-realistic 3D head avatars, stylized 3D avatars emphasize the expression of personal identity and the faithful transfer of target styles.
Blendshape is an efficient facial animation representation method that can generate continuous and high-quality expressions through a set of 3D meshes, each corresponding to a specific facial expression. These facial shapes are generated by linearly blending the basis meshes using weighting coefficients. With the advent of Neural Radiance Fields (NeRF) (Mildenhall et al., 2021), Gao et al. (Gao et al., 2022) and Zheng et al. (Zheng et al., 2022) introduced the blendshape concept into NeRF, enabling avatar animation through the construction of a group of NeRF blendshapes that are linearly blended. Furthermore, the recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) significantly improves rendering efficiency while achieving high-quality head reconstruction compared to NeRF. Based on this, 3D Gaussian Blendshapes (3DGB) (Ma et al., 2024) successfully integrated the blendshape concept with Gaussian splatting, achieving both fast rendering speeds and state-of-the-art performance in head reconstruction and animation.
In contrast to the aforementioned studies that focus on reconstructing photo-realistic head avatars, we propose ToonifyGB, a two-stage framework for stylized 3D head avatar synthesis and animation. Given monocular input video frames, Stage 1 employs an improved StyleGAN to generate a more stable (i.e., less jittery) stylized video beyond pre‑aligned cropping. In Stage 2, we begin with 3DGB to reconstruct a neutral head model with a set of expression blendshapes, each represented as 3D Gaussians, as shown in 1. Finally, by incorporating the facial tracker (Zielonka et al., 2022), we use the tracked motion parameters to drive ToonifyGB in synthesizing the animated head avatar.
In summary, the contributions of this work are as follows:
  • We propose ToonifyGB, an efficient two-stage framework for synthesizing 3D stylized head avatars from input videos using Gaussian blendshapes. This framework supports diverse toonified styles and enables real-time animation.
  • We demonstrate that reducing frame-by-frame jitter enables Gaussian blendshapes to better capture the high-frequency details in the generated video, improving the quality of 3D stylized head avatar animations.
  • To the best of our knowledge, we are the first to synthesize 3D stylized head avatars using Gaussian blendshapes.
2. Related Work
2.1. StyleGAN and Toonify
StyleGAN (Karras et al., 2019, 2020) has been widely used to generate realistic facial images across diverse styles. Inversion of StyleGAN enables projecting real facial images into its latent space, allowing for subsequent edits such as adding glasses or changing hairstyle or age (Abdal et al., 2019; Patashnik et al., 2021). To enhance inversion efficiency, methods such as pSp (Richardson et al., 2021) and e4e (Tov et al., 2021) employ encoders to directly project target faces into their corresponding latent codes. However, these methods often struggle to reconstruct fine image details, resulting in unsatisfactory reconstruction quality. To address these limitations, ReStyle (Alaluf et al., 2021) and HFGI (Wang et al., 2022) improve reconstruction fidelity by respectively predicting latent code residuals and correcting middle layer features. Nevertheless, these methods remain limited to aligned and cropped facial images to achieve effective editing and reconstruction.
Recently, researchers (Pinkney and Adler, 2020; Ojha et al., 2021; Jang et al., 2021; Yang et al., 2022a; Gal et al., 2022) have explored using StyleGAN for target‑domain image generation through transfer learning. Among these works, Toonify (Pinkney and Adler, 2020) fine‑tunes the trained generator to blend realistic textures with toonified facial structures. In addition to image editing, StyleGAN has also been widely applied to video editing. Related studies have focused on enhancing video editing performance by employing temporal correlations in low‑dimensional latent codes (Fox et al., 2021), disentangling identity from facial attributes (Yao et al., 2021), incorporating sketch‑based branches (Liu et al., 2022), and tuning the generator to maintain temporal consistency (Tzaban et al., 2022). However, these methods typically rely on facial alignment and cropping as preprocessing steps. Although StyleGAN3 (Karras et al., 2021) was introduced to support unaligned face inputs, a subsequent study (Alaluf et al., 2022) has shown that it struggles to encode facial features effectively without preprocessing, often resulting in structural artifacts. To overcome these limitations, methods such as VToonify (Yang et al., 2022b) and StyleGANEX (Yang et al., 2023) have been proposed to directly process videos beyond pre-aligned cropping. Nevertheless, these methods remain limited to 2D image representations and have yet to be extended to 3D applications.
Figure 2. Pipeline: Our ToonifyGB framework consists of two stages: Stage 1 involves the generation of stylized videos, and Stage 2 focuses on the synthesis of 3D stylized head avatars using Gaussian blendshapes.
[…]
The architecture of our method ToonifyGB.
2.2. 3D Head Avatar
Since the introduction of NeRF (Mildenhall et al., 2021), implicit representation-based methods (Yenamandra et al., 2021; Zheng et al., 2022; Hong et al., 2022; Chan et al., 2022; Xu et al., 2023) for head reconstruction have achieved remarkable progress. 3DGS (Kerbl et al., 2023) has obtained a significant breakthrough in 3D reconstruction, further advancing the development of downstream applications such as 3D head modeling. Although several Gaussian-based methods (Qian et al., 2024; Xu et al., 2024; Chen et al., 2024; Ma et al., 2024; Xiang et al., 2024; Abdal et al., 2024; Kirschstein et al., 2024) have demonstrated high-quality head reconstruction and impressive rendering performance, they typically focus on photo-realistic avatars, with relatively limited exploration of avatar stylization. Stylized head avatars, characterized by geometric abstraction and artistic expression, differ significantly from the photo-realistic avatars synthesized by the aforementioned methods.
Pre-trained 3D GANs (Wu et al., 2016) enable high-quality generation, making 3D head stylization possible. Although fine-tuning 3D generators for geometric and texture-based stylization has proven effective (Or-El et al., 2022; Jin et al., 2022; Abdal et al., 2023; Lan et al., 2023), performing independent fine-tuning for each new style remains costly. Toonify3D (Jang et al., 2024) addressed this limitation by predicting facial surface normals using the proposed StyleNormal, enabling direct face stylization without additional fine-tuning. Similarly, DeformToon3D (Zhang et al., 2023) introduced StyleField to predict conditional 3D deformations, aligning NeRF representations in real space with style space to achieve geometric stylization and obviate per-style fine‑tuning. However, Toonify3D suffers from limited data diversity, and DeformToon3D cannot support novel-view animations, which limits their application scenarios.
3. Method
3.1. ToonifyGB Framework
Given a monocular video input, ToonifyGB applies frame-by-frame stylization to generate the corresponding stylized frames. For inputs such as live streams or selfie videos, the face often occupies only a small portion of each frame, while the rest includes the hairstyle and upper body. Traditional methods (Karras et al., 2019, 2020) typically require facial alignment, cropping, and editing before synthesizing the results back into the original frame. This process often introduces visual discontinuities at the seams, resulting in noticeable jitter in the output video. To address this problem, we employ an improved StyleGAN based on StyleGANEX (Yang et al., 2023) in Stage 1, enabling stable stylized video generation at the original resolution, as shown in Figure 2.
To prepare the training data for Stage 2, we follow (Zielonka et al., 2023; Ma et al., 2024), using the facial tracker from (Zielonka et al., 2022) to compute FLAME (Li et al., 2017) meshes for a neutral head model and a set of basis expression blendshapes. This process also provides camera parameters $C$, joint and pose parameters $\Theta$, and expression coefficients $\{\psi_{k}\}$ for each frame. In addition to enabling facial expressions control, the FLAME model based on Principal Component Analysis (PCA) provides joint and pose parameters for controlling head, eyeball, eyelid, and jaw movements. As shown in Figure 2, we apply Linear Blend Skinning (LBS) to transform 3D head avatar: $B^{\psi*}=LBS(B^{\psi},\Theta)$. The transformed Gaussian model is then rendered into a Gaussian head avatar in real-time using Gaussian Splatting. Finally, by incorporating camera parameters, we enable head avatar animation from novel views.
Figure 3. Generators of StyleGAN2 and ours: For clarity and ease of comparison, we omit components that are exactly the same: the learned weights, biases, and noises.
[…]
StyleGAN2 generator vs. our generator.
3.2. Stylized Video Generation
As demonstrated by StyleGANEX (Yang et al., 2023), the fixed $4\times 4$ resolution of the first‑layer feature map in StyleGAN2 (Karras et al., 2020) is insufficient for capturing fine details in non‑aligned faces. In contrast, the seventh layer produces a higher‑resolution $32\times 32$ feature map, enabling more accurate facial generation. Based on this, we propose enhancing the feature maps of the seven shallow layers in StyleGAN2 to the $32\times 32$ resolution. As shown in Figure 3, our improvements to StyleGAN2 are as follows. First, we replace the constant $4\times 4$ input of the first layer with a variable feature map of resolution $H/32\times W/32$, thus supporting arbitrary input sizes. Then, we substitute the standard convolutions in the shallow layers with dilated convolutions to enlarge the receptive field. Finally, we remove all upsample operations before the eighth layer, ensuring that the seven shallow layers maintain the same $32\times 32$ resolution. These improvements effectively address the limitations beyond pre-aligned cropping. Our improved StyleGAN can stylize input videos without requiring facial alignment and cropping as preprocessing, as shown in Figure 9.
3.3. Gaussian Blendshapes Synthesis
We represent all Gaussian head avatars using 3D Gaussians. As described in 3DGS (Kerbl et al., 2023), each Gaussian has some basic properties including Gaussian center $\mu$, scale $s$, color $c$, opacity $\alpha$, and rotation $q$. Based on 3DGB (Ma et al., 2024), our Gaussian blendshape representation consists of a neutral model $B_{0}$ and a set of $n$ expression blendshapes ${B_{1},B_{2},\ldots,B_{n}}$. Each Gaussian in the neutral model $B_{0}$ has a set of blend weights $w$ to control joint and pose. In addition, each Gaussian in an expression blendshape $B_{k}$ corresponds one‑to‑one to a Gaussian in the neutral model $B_{0}$. The difference between $B_{k}$ and $B_{0}$ is defined as the difference in their corresponding Gaussian properties: $\Delta B_{k}=B_{k}-B_{0}$. The expression of Gaussian head avatar $B^{\psi}$ can be computed as follows:
(1) $B^{\psi}=B_{0}+\sum_{k=1}^{n}\psi_{k}\,\Delta B_{k}$
where $\psi_{k}$ are the expression coefficients.
Since the FLAME meshes and blendshape models do not include interior mouth components such as teeth, we adopt the method of 3DGB (Ma et al., 2024) by defining a separate set of Gaussians for the mouth $B_{m}$. The properties of these mouth Gaussians are not affected by expression changes, they only move with the jaw joint in the FLAME model. The mouth for the Gaussian head avatar $B_{m}^{*}$ is computed as:
(2) $B_{m}^{*}=\mathrm{LBS}(B_{m},\Theta),$
where $\mathrm{LBS}$ denotes linear blend skinning with the joint and pose parameters $\Theta$. In summary, both $B^{\psi*}$ and $B_{m}^{*}$ are rendered together to form the complete Gaussian head avatar, as shown in Figure 2.
3.4. Loss Function
We adopt the loss function from 3DGB (Ma et al., 2024), and define the total loss as follows:
(3) $L=\lambda_{1}L_{rgb}+\lambda_{2}L_{\alpha}+\lambda_{3}L_{reg},$
where the default values of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are 1, 10, 100, respectively.
$L_{rgb}$ encourages the rendered image to match the video frame in both color and structure. It is computed as a weighted blend of an $L_{1}$ loss and a D-SSIM loss:
(4) $L_{rgb}=\lambda_{rgb}L_{1}+(1-\lambda_{rgb})L_{D-SSIM},$
where the default value of $\lambda_{rgb}$ is 0.2.
$L_{\alpha}$ penalizes opacity outside the head mask. For each frame $i$, we compute the accumulated opacity image $I_{\alpha}$ and head mask $M_{h}^{i}$, then average over $F$ frames:
(5) $L_{\alpha}=\frac{1}{F}\sum_{i=1}^{F}\big{\|}I_{\alpha}^{i}-M_{h}^{i}\big{\|}_{2},$
$L_{reg}$ keeps mouth Gaussians within a predefined cylindrical volume $V$. Let $\{\mathbf{x}_{i}\}_{i=1}^{N}$ be the Gaussian centers inside the mouth. we penalize any point outside $V$ using the signed distance function $SDF(\mathbf{x}_{i},V)$, the equation of $L_{reg}$ loss is as follows:
(6) $L_{reg}=\frac{1}{N}\sum_{i=1}^{N}\left(\max\left(SDF(\mathbf{x}_{i},V),0\right)\right)^{2}$
where $N$ is the number of mouth Gaussians.
4. Experiments
4.1. Baseline and Dataset
4.1.1. Baseline
Due to the current lack of methods for synthesizing 3D stylized head avatars using Gaussian blendshapes, we compare our method with the following state‑of‑the‑art methods for synthesizing photo‑realistic 3D head avatars: INSTA (Zielonka et al., 2023), PointAvatar (Zheng et al., 2023), FLARE (Bharadwaj et al., 2023), SplattingAvatar (Shao et al., 2024), FlashAvatar (Xiang et al., 2024), and 3DGB (Ma et al., 2024). Notably, 3DGB shares the similar architecture with ours but focuses on photo‑realistic avatar synthesis, and does not support the synthesis of diverse stylized avatars.
4.1.2. Dataset
We evaluate state-of-the-art methods for photo-realistic avatar synthesis and our method using six videos from INSTA (Zielonka et al., 2023) dataset. Each video is cropped and resized to $512\times 512$ resolution, with sequence lengths ranging from 1,000 to 4,000 frames.
Following the method of 3DGB (Ma et al., 2024), we retain the final 350 frames of each video for testing. Both methods (3DGB (Ma et al., 2024) and ours) apply the same preprocessing techniques (Zielonka et al., 2022, 2023), including background removal and FLAME parameter extraction.
4.2. Evaluation Metrics
We employ two metrics to evaluate video stabilization: Inter-frame Transformation Fidelity (ITF) (Morimoto and Chellappa, 1998; Marcenaro et al., 2001; Xu et al., 2012) and Inter-frame Similarity Index (ISI) (Guilluy et al., 2018; James et al., 2023). ITF measures the inter-frame Peak Signal-to-Noise Ratio (PSNR) in dB based on the mean squared error. The intuitive idea of ITF is that a more stable video (i.e., less jittery) will have greater similarity between adjacent frames compared to an unstable version of the same video. ISI computes the average Structural Similarity (SSIM) between adjacent frames across the video. Higher ISI values indicate greater perceptual similarity between frames, leading to improved visual comfort for viewers.
For 3D head avatar synthesis, we evaluate the performance of state-of-the-art methods for photo-realistic avatar synthesis and our method using standard evaluation metrics (Zhang et al., 2018; Ma et al., 2024), including PSNR (in dB), and SSIM. In addition, we record the training time for each result in minutes and the rendering speed in frames per second (fps).
4.3. Implementation Details
To ensure a fair performance comparison, the training and testing of all methods are performed on a single RTX 4090 GPU. Our methods are implemented in Python using the PyTorch framework. For 2D stylized video generation, we use the pre-trained models provided by StyleGANEX (Yang et al., 2023). For training the 3D stylized head avatars, we employ the Adam optimizer (Kingma and Ba, 2015), setting the initial learning rates of the Gaussian properties $\{\mathbf{x}_{k},\alpha_{k},\mathbf{s}_{k},\mathbf{q}_{k},SH_{k}\}$ to $3.2\times 10^{-7}$, $5\times 10^{-5}$, $5\times 10^{-4}$, $1\times 10^{-4}$, and $1.25\times 10^{-3}$, respectively. Following 3DGB (Ma et al., 2024), the initial number of sampled Gaussians is 50k for the neutral head model and 14k for the mouth interior.
Figure 4. Visualization of stylized video generation: We present details of the real head from the input video, and the “Arcane” stylized head generated by our method. From left to right, the results for the data samples “bala” and “wojtek_1” are shown. The generated results are highly reproducible, preserving key features of the original head, such as the beard, mouth shape, and the black mole above the eye.
[…]
First row focuses on sides, second row on details.
Table 1. Duration (in seconds) of the input videos, and inference time (in seconds) of our method. The input videos are $512\times 512$ pixels. The inference is done on an RTX 4090 GPU.
Dataset justin malte_1 nf_01 bala wojtek_1 person_0004
Duration 98 130 130 159 137 60
Inference 221 260 213 342 275 108
Table 2. Quantitative comparison of Video Stabilization: We compare the original input (OI), the aligned input (AI), our “Arcane” (OA), and the aligned “Arcane” (AA) videos.
Dataset justin malte_1 nf_01 bala wojtek_1 person_0004
[…]
AI 32.45 28.84 26.49 27.97 29.09 34.80
OA 35.80 34.51 29.36 36.01 36.77 33.82
AA 31.35 26.04 25.31 26.43 28.31 30.84
[…]
AI 0.9066 0.9276 0.8918 0.8995 0.9126 0.9270
OA 0.9700 0.9643 0.9382 0.9685 0.9670 0.9532
AA 0.9034 0.8965 0.8887 0.8963 0.9030 0.9036
4.4. Quantitative Comparison
4.4.1. Video Stabilization
We employ an improved StyleGAN model to generate six videos in the “Arcane” style. The duration of each video and the corresponding inference time are presented in Table 1. For videos lasting between 60 and 160 seconds, the generation inference time ranges from approximately 100 to 350 seconds.
We adopt a common facial alignment method using the facial keypoint predictor (Kazemi and Sullivan, 2014) for input videos. We compare the original input videos (OI) with the aligned input videos (AI), and compare the “Arcane” style videos generated by our method (OA) with those generated from AI (AA). As shown in Table 2, both IFT and ISI values of AI are consistently lower than those of OI. Similarly, AA obtain lower IFT and ISI values than OA. These results indicate that applying facial alignment and cropping before frame-by-frame generation (i.e., AI and AA) produces more unstable (i.e., more jittery) videos.
Table 3. Quantitative comparison of 3D head avatars: We evaluate our method and state‑of‑the‑art methods on the INSTA (Zielonka et al., 2023) dataset. In each metric group, the best value is highlighted in bold, and the second‑best is underlined.
Method justin malte_1 nf_01 bala wojtek_1 person_0004
[…]
INSTA 31.66 0.9591 27.44 0.9159 26.45 0.8937 29.53 0.8896 31.36 0.9452 25.44 0.8478
PointAvatar 30.40 0.9373 24.98 0.8853 25.25 0.8919 27.88 0.8658 28.82 0.9192 23.29 0.8576
FLARE 29.10 0.9363 25.93 0.8973 25.97 0.9027 27.20 0.8761 27.84 0.9216 25.53 0.9015
SplattingAvatar 30.93 0.9482 27.66 0.9243 27.08 0.9202 32.14 0.9272 29.54 0.9400 26.49 0.9075
FlashAvatar 32.16 0.9611 27.45 0.9326 28.02 0.9326 30.27 0.8494 32.02 0.9509 25.49 0.8996
3DGB 32.63 0.9643 28.65 0.9432 28.06 0.9340 33.29 0.9457 32.57 0.9623 23.66 0.8449
Ours (Arcane) 33.12 0.9628 29.55 0.9360 28.33 0.9288 33.39 0.9488 30.56 0.9436 28.76 0.9110
Ours (Pixar) 33.42 0.9662 27.01 0.9375 28.34 0.9341 30.84 0.9337 31.14 0.9583 23.16 0.8338
Table 4. Performance comparison: We record the training time (in minutes) and the rendering speed (in fps) of 3DGB and our method in both “Arcane” (A) and “Pixar” (P) styles.
Method Metric justin malte_1 nf_01 bala wojtek_1 person_0004
[…]
Ours (A) 40 45 44 45 50 44
Ours (P) 43 40 43 44 45 44
[…]
Ours (A) 140 142 131 135 140 128
Ours (P) 141 133 128 132 134 127
4.4.2. 3D Head Avatar
We evaluate our method and state‑of‑the‑art methods using standard metrics for animatable head reconstruction, as well as training and rendering times. The quantitative results are presented in Table 3, and the training and rendering times are recorded in Table 4. With the additional integration of stylization, our method achieves performance comparable to the state‑of‑the‑art on the PSNR and SSIM metrics in most cases, and even outperforms them on certain data. Specifically, our method outperforms all other methods on synthesizing the “Arcane” style for the “bala” and “person_0004” data, as well as the “Pixar” style for the “justin” and “nf_01” data.
In addition, although our method integrates stylization into 3D head avatars, its training and rendering times remain comparable to those of the method of 3DGB. In certain cases, our method is even more efficient in both training and rendering. Combined with the additional time required for video generation (as shown in Table 1), the overall time cost of our method remains acceptable.
Figure 5. Qualitative comparison of each stage: We present the input video head frames, the corresponding stylized videos, and the 3D head avatars synthesized by our method.
[…]
Qualitative comparison between video and 3D head avatars.
Figure 6. Qualitative comparison of 3D Gaussian head avatars: We present 3D head avatars using Gaussian blendshapes synthesized by 3DGB (Ma et al., 2024) and our method.
[…]
Qualitative comparison between 3DGB and ours.
4.5. Qualitative Comparison
We present the original video head frames, their corresponding stylized versions generated by our method, and the 3D stylized head avatars synthesized using Gaussian blendshapes. The qualitative comparison is shown in Figure 5. The examples are selected from the “bala” data in the “Pixar” style and the “wojtek_1” data in the “Arcane” style. In the stylized video, the “bala” data exhibits artifacts along the edge of the side of the head. We attribute this to the latent space distribution learned by StyleGAN, which tends to produce striped artifacts when the viewing angle falls outside the distribution covered by the training data. Notably, such artifacts are not present in the corresponding 3D stylized head avatars during rendering. Furthermore, the 3D stylized head avatars successfully preserve fine details from the stylized videos, such as the mole near the eye in the “wojtek_1” data. However, since the 3D avatar synthesis mainly focuses on the facial region, the neck area is typically blurred, as observed in both examples. This blurring also leads to the lower quantitative performance, as the neck region is included in the evaluation.
The qualitative comparison with 3DGB (Ma et al., 2024) is presented in Figures 6 and 11. Our method effectively captures and preserves high-frequency details in the stylized videos. Compared to the method of 3DGB, ToonifyGB can synthesize 3D stylized head avatars with comparable quality and detail.
4.6. Visualization
4.6.1. Generated Video Details
To better demonstrate the visual quality of our generated videos, we present several examples in Figure 9, and select two representative videos for detailed comparison in Figure 4. Specifically, we show real head frames from the “bala” and “wojtek_1” videos, as well as the corresponding heads of generated videos in the “Arcane” style. The results demonstrate that key facial features, such as the beard, mouth shape, and even small details like the black mole above the eye in the lower right image, are well preserved after the stylization process. These details highlight the excellent performance of our method in terms of detail preservation and identity consistency.
4.6.2. Gaussian Blendshapes
We present eight Gaussian blendshapes for the “justin” sample, as well as their corresponding mesh blendshapes, as shown in Figure 10. For more visualizations, please refer to the video in our supplementary material.
Table 5. Ablation Study: We compare 3D head avatars (using “justin” data generated in both styles) synthesized from different input videos: one generated by our method, and the other using facial alignment and cropping as prprocessing.
[…]
w/ Face Align & Crop 32.23 0.9387 0.1587
Ours 33.27 0.9645 0.0796
Figure 7. Ablation Study: We present the 3D head avatar synthesized by our method, as well as that synthesized using the cross-identity reenactment from 3DGB (Ma et al., 2024).
[…]
Ablation study on our method and cross-identity reenactment from 3DGB.
4.7. Ablation Study
4.7.1. Facial Alignment and Cropping
We compare 3D stylized head avatars synthesized from videos stylized by our pipeline with those synthesized from videos preprocessed by facial alignment and cropping. We evaluate the resulting avatars using PSNR, SSIM, and Learned Perceptual Image Patch Similarity (LPIPS) metrics. The results are shown in Table 5. The avatars synthesized by our method outperform those synthesized from videos with facial alignment and cropping as preprocessing in all evaluation metrics.
4.7.2. Cross-identity Reenactment
One of the key contributions of this work is the real-time synthesis of stylized head avatar animations using expression blendshapes for each frame. The method of 3DGB (Ma et al., 2024) supports cross-identity reenactment, enabling the transfer of facial expressions between different head avatars while preserving the target’s personal attributes. However, transferring the expressions from a real head to a rendered stylized head avatar, as shown in Figure 7, may produce unsatisfactory results, particularly around the mouth region. This error occurs because the source heads are stylized, and the expression blendshapes differ significantly. This demonstrates the importance of video-based acquisition, as 3DGB method based on cross-identity reenactment is insufficient for efficient animation synthesis.
Figure 8. Limitation: We present side-view renderings synthesized by 3DGB and our method.
[…]
Side View of 3DGB and ours on “justin” data.
5. Limitation and Future Work
Our method struggles to render side views of 3D stylized head avatars if the training data (i.e., input video) lacks side views of the real head. As shown in Figure 8, this limitation also affects the baseline method. In fact, existing NeRF-based and Gaussian-based methods have yet to effectively address this problem. Rendering novel views from single view training data remains an open problem for future work. Employing 2D GANs to synthesize videos with side views as additional training data and enhancing the generalization ability of our model are two potential solutions, which are interesting directions for future investigation.
6. Conclusion
We propose a novel two-stage framework, ToonifyGB, which uses Gaussian blendshapes to synthesize the head animations in diverse toonified styles from input videos. We incorporate an improved StyleGAN model to generate stylized videos beyond facial alignment and cropping as preprocessing in Stage 1. This process produces more stable video outputs, providing a basis for high-quality 3D head avatar animation synthesis. ToonifyGB serves as a bridge between 2D generation and 3D modeling, offering valuable insights into applying 2D GANs for 3D head reconstruction.